Grade 1210 grade maths

Without actually performing the long division, state whether the following rational numbers will have a terminating decimal expansion or a non-terminating repeating decimal expansion:
17/8

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Factorizing the denominator, we get,
8 = 2×2×2 = 2^3
Since, the denominator has only 2 as its factor,17/8has a terminating decimal expansion.

Thanks